OpenStudy (anonymous):

Differentiate: xcos^-1(x) - sqrt(1-x^2)

OpenStudy (freckles):

What troubles are you having on this problem?

OpenStudy (anonymous):

i think im messing up on the multiplication rule on the left side with the x*cos^-1(x). I'm not very familiar with arccos

OpenStudy (freckles):

\[\frac{d \arccos(x)}{dx}=\frac{-1}{\sqrt{1-x^2}}\]